In each of the following scenarios sketch a single spinner with the given characteristics.
A. The outcomes M, A, T and H are each equally likely.
B. The outcomes are R,A,and T with P(R)=3/4, P(A)=1/8 and P(T)=1/8

Found 2 solutions by ikleyn, math_tutor2020:
Answer by ikleyn(52803)

.

Use a pen and a paper.



Answer by math_tutor2020(3817)


Part A

Start by drawing a circle of any size you want.
Then cut it in half, and cut it in half again to get 4 equal slices.


P(M) = 1/4
P(A) = 1/4
P(T) = 1/4
P(H) = 1/4
Add up those probabilities to get 1.
1/4 = 0.25 = 25%

The order of the letters does not matter.


--------------------------------------------------

Part B
Start with the previous drawing. Erase the labels.


Then erase the bottom vertical line and the right-most horizontal line

So far we have a 1/4 slice in the upper left corner.
The remaining chunk is 3/4 of the pizza.

Split the 1/4 slice in half to get two copies of 1/8
(1/4)*(1/2) = 1/8


Then be sure to label the appropriate regions.

P(R) = 3/4 = 0.75
P(A) = 1/8 = 0.125
P(T) = 1/8 = 0.125
Add up those probabilities to get 1.
The order of the "A" and "T" does not matter.

Side note: 3/4 = 6/8